# The Non-Gaussian to Gaussian Transition: Pointwise Heat Kernel Estimates and Optimal Convergence Rates

### 15 Mar 2026 v1

@@ Abstract.
We establish uniform pointwise estimates for the densities of a family of $\alpha$-stable processes with respect to the stability index $\alpha\in[\alpha_{0},2]$ for some $\alpha_{0}>0$, and we estimate the difference between the heat kernels of non-local and local operators, showing that it is controlled by the rate $2-\alpha$. Both estimates (see Proposition 1.5) are new to the literature. Furthermore, as an application, we achieve the optimal rate $2-\alpha$ for the pointwise estimate between the transition probabilities, as well as for the (weighted) total variation and Kantorovich distances between the invariant measures, of anomalous and normal diffusion with drifts that are low in regularity and possibly unbounded. The results on transition probabilities (see Theorem 1.3) are novel, while those on invariant measures (see Theorem 1.8) significantly extend the existing literature.
Keywords: $\alpha$-stable processes; heat kernel estimates; invariant measures; weighted total variation; optimal convergence rates
@@ 1. Introduction
A $\mathbb{R}^{d}$-valued càdlàg process $(L_{t})_{t\geq 0}$ with $d\geq 1$ is called a Lévy process if $L_{0}=0$ almost surely and $L$ has independent and stationary increments. Let $(L_{t}^{(\alpha)})_{t\geq 0}$ be a $d$-dimensional rotationally invariant $\alpha$-stable Lévy process with $\alpha\in(0,2]$, whose characteristic exponent is given by
$\mathbb{E}\mathrm{e}^{i\xi\cdot L_{t}^{(\alpha)}}=\mathrm{e}^{-t|\xi|^{\alpha}}.$ (1.1)
In particular, when $\alpha=2$, $(\frac{1}{\sqrt{2}}L_{t}^{(2)})_{t\geq 0}$ is exactly the standard Brownian motion; when $\alpha\in(0,2)$, it is the so-called standard (or strict) $\alpha$-stable process (cf. [Sa99]). In this paper, we investigate the “distance” (in some suitable sense) between the following two stochastic differential equations (SDEs) driven by $\alpha$-stable processes with varying stability indices:
@@
${\mathord{{\mathrm{d}}}}X_{t}=b(X_{t}){\mathord{{\mathrm{d}}}}t+{\mathord{{\mathrm{d}}}}L^{(2)}_{t}$, (1.3)
where the drift coefficient $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a Borel measurable function.
@@ 1.1. Heat kernel estimates: from non-local to local
It is well known that when $\alpha$ is close to $2$, an $\alpha$-stable process behaves similarly to a Brownian motion, which is clearly observable in computer simulations of sample paths (see, e.g., [JW94]). Roughly speaking, as $\alpha$ approaches $2$, the motion of an $\alpha$-stable process is dominated by small jumps. This behavior is also evident from the perspective of probability densities. Indeed, by (1.1) and Proposition 28.1 of [Sa99], for $\alpha\in(0,2]$, the process $L_{t}^{(\alpha)}$ admits a smooth density function $p^{(\alpha)}(t,\cdot)$ given by the inverse Fourier transform
@@
and all partial derivatives of $p^{(\alpha)}(t,\cdot)$ decay to 0 as $|x|\to\infty$. Moreover, by the dominated convergence theorem, we obtain
@@
Regarding the first question, to the best of our knowledge, no result is currently available in the literature. In present work, we address this gap by establishing a pointwise estimate for the difference between the heat kernels, which yields a convergence rate of $2-\alpha$ (see Proposition 1.4(ii)). A key ingredient in our proof is a novel uniform estimate for the heat kernel with respect to the stability index $\alpha$ (see Proposition 1.4(i)), which is of independent interest.
@@
For the second question, several results concerning the convergence of distributions have already been established. For example, the first named author, Liu, studied in [Liu22a] the convergence behavior, as $\alpha\to 2$, of solutions to SDEs driven by $\alpha$-stable processes with drift belonging to the usual Hölder space. As for quantitative results, we first refer to [LFL23], in which the authors gave a look at the optimal weak convergence rate, showing it to be $2-\alpha$ for drifts that are dissipative, locally bounded, and of Hölder regularity $2+\beta$, by constructing an evolution equation describing the weak difference. Subsequently, Deng et al. [DSX23] established the optimal convergence rate $2-\alpha$ for the Wasserstein-1 distance between $\operatorname{Law}(X_{t}^{(\alpha)}(x))$ and $\operatorname{Law}(X_{t}^{(2)}(y))$. Their proof uses Malliavin calculus and assumes the drift to be dissipative and third-order differentiable. This result was later extended to the case of multiplicative noises in the master’s thesis [Ly24] (written in Chinese), supervised by the first named author, Liu. As a continuation of the work in [DSX23], Deng et al. [DLSX26] derived an estimate for the total variation distance and further extended it to the Wasserstein-$p$ distance for $0<p<1$.
@@
However, a quantitative result on the pointwise convergence of the densities has remained open, even under smooth coefficient assumptions. In this work, we fill this gap by investigating the convergence of heat kernels from SDE (1.2) to SDE (1.3), where the drifts are of low regularity and possibly unbounded, satisfying condition ($\mathbf{H}_{b}^{\beta}$). Our main contribution (Theorem 1.3(ii)) is to establish that, even in this more challenging setting, the convergence rate remains $2-\alpha$. The proof of this result relies crucially on the uniform heat kernel estimates with respect to $\alpha$ established in Theorem 1.3(i). This uniform estimate is, to the best of our knowledge, new in the literature and is of independent interest.
1.1.1. Assumptions and deterministic regularized flows
@@ 1.1.2. Main results
$\Theta:=(d,\kappa_{0},\beta).$
@@
Here, $\theta^{(1)}_{s,t}$, $\varrho_{\alpha}$, and $\phi^{(\eta)}_{\gamma_{1},\gamma_{2}}$ are defined by (1.6), (2.17), and (4.1), respectively.
As a special case (see Lemma 2.5 and Theorem 3.7), we have the following similar estimates for the transition densities of $\alpha$-stable processes themselves.
@@
(i) Assume that $\alpha\in[\alpha_{1},2]$ with some $\alpha_{1}\in(0,2)$, and $j\in{\mathbb{N}}_{0}$. Then there is a constant $c>0$ only depends on $d,\alpha_{1},j$ such that for any $t>0$,
$\nabla^{j}p^{(\alpha)}(t,x)\leq c\varrho_{\alpha}(t,x).$
(ii) Assume that $\alpha\in[\alpha_{2},2]$ with some $\alpha_{2}\in(1,2)$, and $j=0,1,2$. Then there is a constant $c>0$ only depends on $d,\alpha_{2},j$ such that for any $t>0$,
$\begin{split}&|\nabla^{j}p^{(\alpha)}(t,x)-\nabla^{j}p^{(2)}(t,x)|\\ \leq c&\,(2-\alpha)(1+|\ln t|)(1+t^{\frac{\alpha-2}{\alpha}})\sum_{k=1}^{j}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}|x|^{2k-j}\varrho^{(2k)}_{\gamma_{1},\gamma_{2}}(t,x).\end{split}$
where $\varrho^{(j)}_{\gamma_{1},\gamma_{2}}$ is defined by (2.16), and we used the convention $\sum_{k=1}^{0}=\sum_{k=j=0}$.
@@
The above result (ii) remains valid for the stochastic integral $\int_{0}^{t}\sigma(r){\mathord{{\mathrm{d}}}}L_{r}^{(\alpha)}$ provided the coefficient $\sigma(r)$ is diagonal (see Theorem 3.7). In contrast, Theorem 3.1 imposes no such diagonal requirement on $\sigma(r)$. This raises the question of whether the diagonal assumption can be dropped.
@@
By Theorem 1.3 and the equation (1.16) of [MPZ21], it is easy to check that, it is easy to check that for any $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(\frac{12}{7},2)$, there exists a constant $c=c(\alpha_{0},T,\Theta)>0$ such that for each $t\in(0,T]$ and $x,y\in{\mathbb{R}}^{d}$,
$\left\|{\rm Law}(X_{t}^{(\alpha)}(x))-{\rm Law}(X_{t}^{(2)}(y))\right\|_{{\mathrm{var}}}\leq ct^{-1/2}|x-y|+(2-\alpha)t^{\tfrac{7\alpha-16}{2\alpha}}.$
Precisely, we have
@@
$\qquad\lesssim t^{-1/2}|x-y|+(2-\alpha)t^{\tfrac{7\alpha-16}{2\alpha}}.$
1.2. $\alpha$-dependence of invariant measures
Having established the continuous dependence of the heat kernels for solutions to SDEs on the stability index $\alpha$ (see Theorem 1.3), we then apply this result to derive, under an additional dissipation condition, the (weighted) total variation and Kantorovich distances for the associated invariant measures (see Theorem 1.7).
There are a few results concerning the convergence of invariant measures for stochastic systems in the transition from non-Gaussian to Gaussian noise. In this direction, we refer the reader to a series of works [Liu22b, Liu22c, LL25], in which the authors investigated the convergence of ergodic measures for three types of stochastic equations driven by cylindrical $\alpha$-stable processes as $\alpha\to 2$: the stochastic real Ginzburg–Landau equation (cf. [Liu22b]), the stochastic Burgers equation (cf. [Liu22c]), and the stochastic Navier–Stokes equation (cf. [LL25]). These works provide a qualitative understanding of the convergence.
References Convergence Method Conditions
1.2.1. Assumptions and distances between measures
@@ 1.2.2. Main results
$\mathcal{T}_{p}(\mu_{1},\mu_{2}):=\inf\left\{{\mathbb{E}}|X-Y|^{p},~\text{law}(X)=\mu_{1},~\text{law}(Y)=\mu_{2}\right\}.$
@@
1.3. Conventions and notations
  • Denote by Beta functions and Gamma functions, respectively,
    and
    Furthermore, notice that
    and
  • For $\alpha\in(0,2)$, define
    $\mathcal{C}(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{2-\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{2-\alpha}{2})}.$ (1.12)
  • In the sequel, we use the conventions $0^{0}=1$ and $\frac{1}{0}=\infty$.
  • For functions $f,g$, define
    and

### 26 Mar 2026 v2

@@ Abstract.
We establish uniform pointwise estimates for the densities of a family of $\alpha$-stable processes with respect to the index $\alpha\in[\alpha_{0},2]$ for some $\alpha_{0}>0$. In addition, we estimate the difference between the heat kernels of non-local and local operators, showing that it is controlled by the rate $2-\alpha$. Both estimates (see Proposition 2.4) are new to the literature. Furthermore, as an application, we achieve the optimal rate $2-\alpha$ for the pointwise estimate between the transition probabilities, as well as for the (weighted) total variation and Kantorovich distances between the invariant measures, of non-Gaussian and Gaussian diffusion. These results are obtained under the assumption that the drifts are locally $\beta$-Hölder continuous, with the latter additionally requiring dissipativity. The results on transition probabilities (see Theorem 2.3) are novel, while those on invariant measures (see Theorem 2.7) significantly extend the existing literature.
Key words and phrases:
$\alpha$-stable processes; heat kernel estimates; invariant measures; weighted total variation; optimal convergence rates
2020 Mathematics Subject Classification:
Primary: 60G52 Stable stochastic processes, 35K08 Heat kernel; Secondary: 60B10 Convergence of probability measures, 60H10 Stochastic ordinary differential equations (aspects of stochastic analysis)
@@ 1. Introduction
1.1. Motivation
A Lévy process $(L_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$ is called $\alpha$-stable if it is self-similar with stability index $\alpha\in(0,2]$; that is, for any $c>0$, the scaled process $(c^{-1/\alpha}L_{ct})_{t\geq 0}$ has the same distribution as $(L_{t})_{t\geq 0}$. Here $\alpha$ is called the stability index (which also serves as the tail index: the smaller $\alpha$, the heavier the tail). When $\alpha=2$, the process is (up to a scaling factor) a standard Brownian motion, which is the only Gaussian Lévy process. When $\alpha\in(0,2)$, the process is recognized as a non-Gaussian process.
It is well known that when $\alpha$ is close to $2$, an $\alpha$-stable process behaves similarly to a Brownian motion, which is clearly observable in computer simulations of sample paths (see, e.g., [JW94]). Roughly speaking, as $\alpha$ approaches $2$, the motion of an $\alpha$-stable process is dominated by small jumps (see also [DZ16, SZ20]). The limit $\alpha\to 2$ thus connects two different types of stochastic models: from non-Gaussian to Gaussian behavior.
In recent years, SDEs driven by $\alpha$-stable processes have been studied across many fields, where the stability index $\alpha$ governs the shift from heavy-tailed, jump-type models to Gaussian continuous models. In physics, this shift characterizes the evolution from anomalous diffusion to the classical equilibrium of Brownian motion. The jumped processes also appear in models of collisions in certain physical systems, such as Boltzmann equation (see e.g., [Ta78]), and the Surface Quasi-Geostrophic system (see e.g., [HRZ24]). In engineering, non-Gaussian $\alpha$-stable distributions are used to characterize signals that exhibit sharp spikes or occasional bursts, where the transition $\alpha\to 2$ corresponds to a return to Gaussian signals (see e.g., [SCQ]). In economics, heavy-tailed distributions—such as Pareto’s law for income and wealth (see e.g., [Ar15])—are important in extreme value theory and risk management (see e.g., [EKM]); $\alpha$-stable processes provide a flexible family for describing such tail behavior. In financial mathematics, the heavy tails and jumps of asset returns are well captured by $\alpha$-stable distributions. A stability index $\alpha<2$ reflects their heavy-tailed nature and “black swan” events (see e.g., [No]), while $\alpha=2$ reduces to the Gaussian case, corresponding to the classical Black-Scholes framework (see e.g., [Ok03]).
Although $\alpha$-stable processes appear in many areas, a quantitative understanding of the limit $\alpha\to 2$ has received little attention in the literature. In this paper, we study the “distance” (in some suitable sense) between the following two stochastic differential equations (SDEs):
@@ 1.1. Motivation
${\mathord{{\mathrm{d}}}}X^{(2)}_{t}=b(X^{(2)}_{t}){\mathord{{\mathrm{d}}}}t+{\mathord{{\mathrm{d}}}}L^{(2)}_{t}$, (1.2)
where $b:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a Borel measurable drift coefficient, and $(L_{t}^{(\alpha)})_{t\geq 0}$ is the rotationally invariant $\alpha$-stable Lévy process with $\alpha\in(0,2]$. In particular, for SDEs with drift coefficients that are of low regularity—a condition that occurs naturally in many applications—the convergence rates as $\alpha\to 2$ are not covered by existing results. In this work, we study this problem using heat kernel estimates. We obtain convergence rates for the densities of solutions and invariant measures to the corresponding SDEs. To this end, we establish uniform heat kernel estimates with respect to the stability index $\alpha$. The main results are presented in Theorems 2.3 and 2.7.
@@ 1.2. Heat kernel estimates: from non-local to local
As a matter of fact, for rotationally invariant $\alpha$-stable Lévy processes $L^{(\alpha)}$, the limit $\alpha\to 2$ is evident when one looks at the probability densities. Notice that the characteristic function (see e.g., [Sa99]) is given by
$\mathbb{E}\mathrm{e}^{i\xi\cdot L_{t}^{(\alpha)}}=\mathrm{e}^{-t|\xi|^{\alpha}},$ (1.3)
which, by Proposition 28.1 of [Sa99], derives that $L_{t}^{(\alpha)}$ admits a smooth density function $p^{(\alpha)}(t,\cdot)$ given by the inverse Fourier transform
@@
and all partial derivatives of $p^{(\alpha)}(t,\cdot)$ decay to 0 as $|x|\to\infty$. When $\alpha=2$, the process reduces to Brownian motion, whose density $p^{(2)}(t,\cdot)$ corresponds to the Laplacian (a local generator). When $\alpha<2$, the density $p^{(\alpha)}(t,\cdot)$ is associated with the fractional Laplacian (a non-local generator). By the dominated convergence theorem, we obtain the convergence
@@
Regarding the first question, to the best of our knowledge, no result is currently available in the literature. In present work, we address this gap by establishing a pointwise estimate for the difference between the heat kernels, which yields a convergence rate of $2-\alpha$ (see Proposition 2.4(ii)). A key ingredient in our proof is a novel uniform estimate for the heat kernel with respect to the index $\alpha$ (see Lemma 3.5), which is of independent interest. Precisely, for $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$, there exists a constant $c>0$ depending only on $d,\alpha_{0}$ such that for any $t>0$,
$p^{(\alpha)}(t,x)\lesssim_{c}t^{-d/\alpha}\wedge\frac{t}{|x|^{d+\alpha}}=:\varrho_{\alpha}(t,x).$
Compared with classical results (e.g., Lemma 2.2 in [CZ16]), the difference is that our constant $c$ does not depend on the index $\alpha$. See Proposition 2.4(i) for a similar result for the derivatives.
@@
For the second question, several results concerning the convergence of distributions have already been established. For example, the first named author, Liu, studied in [Liu22a] the convergence behavior, as $\alpha\to 2$, of solutions to SDEs driven by $\alpha$-stable processes with drift belonging to the usual Hölder space. As for quantitative results, we first refer to [LFL23], in which the authors gave a look at the optimal weak convergence rate, showing it to be $2-\alpha$ for drifts that are dissipative, locally bounded, and of Hölder regularity $2+\beta$, by constructing an evolution equation describing the weak difference. Subsequently, Deng et al. [DSX23] established the optimal convergence rate $2-\alpha$ for the Wasserstein-1 distance between $\operatorname{Law}(X_{t}^{(\alpha)}(x))$ and $\operatorname{Law}(X_{t}^{(2)}(y))$. Their proof uses Malliavin calculus and assumes the drift to be dissipative and third-order differentiable. This result was later extended to the case of multiplicative noises in the master’s thesis [Ly24], supervised by the first named author, Liu. As a continuation of the work in [DSX23], Deng et al. [DLSX26] derived an estimate for the total variation distance and further extended it to the Wasserstein-$p$ distance for $0<p<1$.
@@
However, a quantitative result on the pointwise convergence of the densities has remained open, even under smooth coefficient assumptions. In this work, we fill this gap by investigating the convergence of heat kernels from non-Gaussian diffusion (see SDE (1.1)) to Gaussian diffusion (see SDE (1.2)), where the drifts are of low regularity and possibly unbounded, satisfying condition ($\mathbf{H}_{b}^{\beta}$). Our first main contribution, Theorem 2.3(ii), is to establish that, even in this more challenging setting, the convergence rate remains $2-\alpha$. The proof of this result relies crucially on the uniform heat kernel estimates with respect to $\alpha$ established in Theorem 2.3(i). This uniform estimate is, to the best of our knowledge, new in the literature and is of independent interest.
1.3. $\alpha$-dependence of invariant measures
Having established the continuous dependence of the heat kernels for solutions to SDEs on the index $\alpha$ (see Theorem 2.3), we then apply this result to derive, under an additional dissipation condition, the (weighted) total variation and Kantorovich distances for the associated invariant measures (see Theorem 2.7).
There are a few results concerning the convergence of invariant measures for stochastic systems in the transition from non-Gaussian to Gaussian noise. In this direction, we refer the reader to a series of works [Liu22b, Liu22c, LL25], in which the authors investigated the convergence of ergodic measures for three types of stochastic equations driven by cylindrical $\alpha$-stable processes as $\alpha\to 2$: the stochastic real Ginzburg-Landau equation (cf. [Liu22b]), the stochastic Burgers equation (cf. [Liu22c]), and the stochastic Navier-Stokes equation (cf. [LL25]). These works provide a qualitative understanding of the convergence.
References Convergence Method Conditions
1.4. Conventions and notations
  • Denote by Beta functions and Gamma functions, respectively,
    and
    Furthermore, notice that
    and
  • For $\alpha\in(0,2)$, define
    $\mathcal{C}(d,\alpha):=\frac{\alpha\Gamma(\frac{d+\alpha}{2})}{2^{1-\alpha}\pi^{\frac{d}{2}}\Gamma(\frac{2-\alpha}{2})}.$ (1.9)
  • In the sequel, we use the conventions $0^{0}=1$ and $\frac{1}{0}=\infty$.
  • For functions $f,g$, define
    and
1.5. Outline of paper
The remainder of this paper is organized as follows. In Section 2, we present the main results of this work. Section 3 introduces some basic concepts and properties of isotropic $\alpha$-stable processes, along with elementary estimates. In Section 4, we establish uniform estimates and $\alpha$-continuity results for heat kernels, which serve as the foundation for the subsequent analysis. In Section 5, we study the densities of solutions to SDEs and prove Theorem 2.3 using the parametrix method, which relies on the estimates from Sections 3 and 4. Finally, building on Theorem 2.3, we prove the optimal convergence rates for invariant measures, namely Theorem 2.7, in Section 6.
2. Main results
2.1. Heat kernel estimates
@@
$\Theta:=(d,\kappa_{0},\beta),$
and introduce
$\varrho^{(k)}_{\gamma_{1},\gamma_{2}}(t,x):=t^{-(d+k)/\gamma_{1}}\wedge\frac{t}{|x|^{d+k+\gamma_{2}}},\ \ \text{for}\ \ \gamma_{1},\gamma_{2}\in(0,2],k\geq 0,$
and
$\phi^{(\eta)}_{\gamma_{1},\gamma_{2}}(s,t;x,y):=(t-s)^{\eta-1}\varrho^{(0)}_{\gamma_{1},\gamma_{2}}(t-s,\theta^{(1)}_{s,t}(x)-y),$
where $\theta^{(1)}_{s,t}$ is defined by (2.1). In particular, when $\gamma_{1}=\gamma_{2}=\alpha$, let
$\varrho^{(k)}_{\alpha}(t,x):=\varrho^{(k)}_{\alpha,\alpha}(t,x),\quad\phi^{(\eta)}_{\alpha}(s,t;x,y):=\phi^{(\eta)}_{\alpha,\alpha}(s,t;x,y);$
if $k=0$, let
$\varrho^{(0)}_{\gamma_{1},\gamma_{2}}(t,x):=\varrho_{\gamma_{1},\gamma_{2}}(t,x).$
@@
As a special case (see Theorem 4.1 and Theorem 4.10), we have the following similar estimates for the transition densities, denoted by $p^{(\alpha)}$, of $\alpha$-stable processes themselves.
@@
(i) Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$, and $j\in{\mathbb{N}}_{0}$. Then there is a constant $c>0$ depending only on $d,\alpha_{0},j$ such that for any $t>0$,
$|\nabla^{j}p^{(\alpha)}(t,x)|\leq c\varrho^{(j)}_{\alpha}(t,x).$
(ii) Fix $T>0$. Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$, and $j=1,2$. Then there is a constant $c>0$ depending only on $d,T,\alpha_{0},j$ such that for any $t\in(0,T]$,
$\begin{split}|p^{(\alpha)}(t,x)-p^{(2)}(t,x)|\leq c(2-\alpha)(1+|\ln t|)(1+t^{\frac{\alpha-2}{\alpha}})\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\varrho_{\gamma_{1},\gamma_{2}}(t,x),\end{split}$ (2.4)
and
$\begin{split}&\qquad|\nabla^{j}p^{(\alpha)}(t,x)-\nabla^{j}p^{(2)}(t,x)|\\ &\leq c(2-\alpha)(1+|\ln t|)(1+t^{\frac{\alpha-2}{\alpha}})\sum_{k=1}^{j}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}|x|^{2k-j}\varrho^{(2k)}_{\gamma_{1},\gamma_{2}}(t,x).\end{split}$
@@
The above results (i), and (2.4) in (ii), remain valid for the stochastic integral $\int_{0}^{t}\sigma(r)dL_{r}^{(\alpha)}$ when the coefficient $\sigma(r)$ is invertible and bounded (see Theorem 4.1 and Theorem 4.10(i)). For $j=1,2$, result (ii) holds under the additional condition that $\sigma(r)$ is diagonal (see Theorem 4.10(ii)). This naturally raises the question of whether the diagonal assumption can be relaxed.
@@
For any $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(\frac{12}{7},2)$, there exists a constant $c=c(\alpha_{0},T,\Theta)>0$ such that for each $t\in(0,T]$ and $x,y\in{\mathbb{R}}^{d}$,
$\left\|{\rm Law}(X_{t}^{(\alpha)}(x))-{\rm Law}(X_{t}^{(2)}(y))\right\|_{{\mathrm{var}}}\leq c\left(t^{-1/2}|x-y|+(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}t^{\tfrac{7\alpha-12}{2\alpha}-\frac{\gamma_{2}(d+\gamma_{1})}{\gamma_{1}(d+\gamma_{2})}}\right),$
where $\|\mu_{1}-\mu_{2}\|_{{\mathrm{var}}}:=\sup_{\|h\|_{\infty}\leq 1}|\mu_{1}(h)-\mu_{2}(h)|$ denotes the usual total variance. Precisely, by Theorem 2.3 and the equation (1.16) of [MPZ21], it is easy to check that
@@
$\qquad\lesssim t^{-1/2}|x-y|+(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}t^{\tfrac{7\alpha-12}{2\alpha}-\frac{\gamma_{2}(d+\gamma_{1})}{\gamma_{1}(d+\gamma_{2})}},$
where we used (3.23) in the last inequality.
$2.$ $\alpha$-dependence of invariant measures
@@
$\mathcal{T}_{p}(\mu_{1},\mu_{2}):=\inf\left\{{\mathbb{E}}|X-Y|^{p},~\mathrm{Law}(X)=\mu_{1},~\mathrm{Law}(Y)=\mu_{2}\right\}.$